OpenStudy (winsteria):

Last month, Steven had $4285.68 that he divided into several bank accounts. He deposited 3/8 into an account to save for college. He deposited 1/4 into another account to save for a car. How much money did he deposit to save for college and a car altogether? PLEASE HELP ME??? ILL GIVE YOU A MEDAL AND FAN PLEASE I NEED HELP ASAP PLEASE

OpenStudy (winsteria):

CAN SOMEONE PLEASE HELP ME PLEASE

OpenStudy (mokeira):

\[\frac{ 3 }{ 8 }\times 4285.68= answer\]

OpenStudy (anonymous):

Actually I think it would be \[5/8\times4285.68\] because it says how much money did he deposit to save for college AND a car together

OpenStudy (winsteria):

CAN YOU GIVE ME THE AMOUNT OF MONEY FOR THE ANSWER PLEASE OR??? LIKE WHATS THE ANSWER THIS IS URGENT PLEASE D:

OpenStudy (mokeira):

ok it will be \[(\frac{ 3 }{ 8 }\times 4285.68) +(\frac{ 1 }{ 4 }\times 4285.68)\]

OpenStudy (anonymous):

I concur.

OpenStudy (winsteria):

but that's still not the answer in all?

OpenStudy (mokeira):

simplify it

OpenStudy (mokeira):

i cant give you a straight up answer

OpenStudy (winsteria):

Why??? D:

OpenStudy (winsteria):

im so confussed like really

OpenStudy (anonymous):

It would be considered cheating because this may be directly copied course material.

OpenStudy (anonymous):

We are here to help you arrive at your answers.

OpenStudy (winsteria):

Oh okay so I got $2678.55, can you at least tell me if that is correct???

OpenStudy (anonymous):

Yes. That is correct.

OpenStudy (winsteria):

THANK YOU SO MUCH GOD BLESS YOU